The index of refraction for water is about 1.3. If a light wave moving through the air strikes a surface of water at an angle of 35 degrees (measured to the normal), what angle will the ray have to the normal when it enters the water? (Assume the n value of air is 1)
Answer #1

formula for refractive index is

n1 sin x1 = n2 sin x2

n for air is 1

1 * sin 35 = 1.3 sin x2 => x2 = 26.18 degres
Answer #3

n1* sin x1 = n2* sinx2

sinx2= 1*sin35/ 1.3

= 0.4412126

x2= 26.18127541
Answer #4

BY snell's law

Sin (angle of incidence) / Sin ( Angle of Refraction) = n2/ n1

Sin 350 / Sin @ = 1.3 / 1

Sin @ = Sin 35 / 1.3

Sin @ = 0.44

@ = 26.180ANS